# ChemReasoner: Heuristic Search over a Large Language Model’s Knowledge Space using Quantum-Chemical Feedback

### v3

@@ Abstract
The discovery of new catalysts is essential for the design of new and more efficient chemical processes in order to transition to a sustainable future. We introduce an AI-guided computational screening framework unifying linguistic reasoning with quantum-chemistry based feedback from 3D atomistic representations. Our approach formulates catalyst discovery as an uncertain environment where an agent actively searches for highly effective catalysts via the iterative combination of large language model (LLM)-derived hypotheses and atomistic graph neural network (GNN)-derived feedback. Identified catalysts in intermediate search steps undergo structural evaluation based on spatial orientation, reaction pathways, and stability. Scoring functions based on adsorption energies and barriers steer the exploration in the LLM’s knowledge space toward energetically favorable, high-efficiency catalysts. We introduce planning methods that automatically guide the exploration without human input, providing competitive performance against expert-enumerated chemical descriptor-based implementations. By integrating language-guided reasoning with computational chemistry feedback, our work pioneers AI-accelerated, trustworthy catalyst discovery.
@@ 1 Introduction
The discovery of new catalysts requires one to identify the optimal combination of chemical descriptors (or properties), which are often only empirically understood. Chemists actively reason to mentally search through reactants, catalysts, and operating conditions that enable more energy-efficient chemical conversions. As proposed by seminal work on descriptor-based catalyst search (Nørskov et al., 2011), linking microscopic surface properties to macroscopic catalytic performance via chemical descriptors is key to accelerating such hypothesis generation.
Large Language Models (LLMs) (Wei et al., 2022; Ouyang et al., 2022; Taylor et al., 2022; Lai et al., 2023; OpenAI, 2023) now bring within reach the realization of such data-driven autonomous search to accelerate scientific discovery. Our work aims to enhance natural language reasoning capabilities with quantum-chemical feedback to discover optimal catalysts for target reactions.
The Challenge However, reasoning about complex catalytic processes requires compositionality across multiple modalities, extending beyond the capabilities of existing language models. These include scientific concepts from literature and property prediction with 3D atomistic configurations. Determining the best catalyst is a multi-step process requiring reasoning about multiple macroscopic properties. The first step involves identifying an optimal set of chemical descriptors (e.g. “resistance to poisoning”, “porosity”) which are relevant to the reaction in question. Formally, given a set of these important descriptors , we want to identify the optimal subset to consider when suggesting new catalysts. This yields possible permutations to reason over. The amount of reasoning required scales very quickly, necessitating an external reasoner: an LLM. Using its knowledge of scientific concepts, the LLM both proposes important properties and selects the best possible catalysts (from all possible catalysts) that have these properties. Pruning this large space of candidate catalysts requires reasoning about the complex microscopic interactions that occur between atomistic structures in 3D space based on macroscopic properties. Further, while simple reactions can be assessed via adsorption energies of 3D chemical structures, complex reactions demand consideration of multi-step reaction pathways and selectivity (Unsleber et al., 2023).
@@
We demonstrate the decisive impact of planning methods in automatically navigating chemical search spaces over SOTA LLM-based implementation. Our LLM-planned approach with zero human input (ChemReasoner-Planner) surpasses search guided by expert-selected chemical descriptors (ChemReasoner-Expert) for two out of three categories in our evaluation benchmark.
@@
Third, and uniquely, we establish domain-grounding of language models via quantum chemical property feedback. We go beyond screening catalysts on adsorption energies alone and propose a methodology to reason in terms of reaction pathways and energy barriers.
Our work pioneers an AI-guided approach to computational catalyst screening and discovery. To facilitate community adoption and advancement of this extremely interdisciplinary and compute-intensive research, we commit to releasing all our datasets. This includes query benchmarks from catalysis experts, multi-modal provenance trails, over 700,000+ atomistic trajectory computations, and validation from density functional theory calculations.
@@ 2.2 LLMs for Chemistry
LLMs for chemistry can be divided into two categories: domain-specific models and adapted general-purpose models. Multimodal, domain-specific molecule-language models have recently emerged to target a number of problems in the chemistry domain (Edwards et al., 2021; Vall et al., 2021; Zeng et al., 2022; Xu & Wang, 2022; Su et al., 2022; Edwards et al., 2022a). For brevity, we discuss them further in Appendix D. On the other hand, general-domain LLMs such as GPT-4 (OpenAI, 2023) have been adopted in an agent-based approach to interface with chemistry-specific tools, allowing information gathering and hypothesis generation (Boiko et al., 2023; Bran et al., 2023). While this work is exciting, it differs from our approach, where we employ an overarching, domain-specific reward function to probe the LLM’s knowledge, enabling stronger, simulation-backed guarantees about complex scientific reasoning. We note that our method builds upon Sprueill et al. (2023). Unlike their approach, which uses LLM-computed rewards, we integrate true computational chemistry-based rewards to guide the model. Further, we integrate a context-aware planner into our search algorithm to automatically guide the search.
3 System and Methods
3.1 Heuristic Search
The goal of heuristic search is to answer a user-specified natural language query by systematically exploring candidates from different regions in chemical space. This is accomplished by taking the original query (or prompt) and corresponding LLM answers, and systematically applying different screening criteria to iteratively contextualize the LLM prompts and answers into a narrowed region of the chemical space. This process is illustrated in Figure 2. In this paper, we employ the beam search method (Rubin & Reddy, 1977) for heuristic search.
@@ 3.1 Heuristic Search
.append
Formally, our goal is to search through chemical descriptors to design the optimal prompt, which leads the LLM to return the best candidate catalysts for a catalysis query. Starting with a general prompt , we use a set of actions to modify the prompt to improve the LLM output with respect to a reward function, . Notably, ChemReasoner-Planner generates its own action space (Figure 3).
Definitions: We define the Search Tree as a hierarchical tree consisting of (prompt, answer, and reward) nodes. Each node in this tree represents a state in the search space (henceforth referred to as search state or query state). Nodes are linked if an action modifies the prompt of one into the other. We denote a path from root to leaf node a Reasoning Pathway.
Following Sprueill et al. (2023), each LLM prompt has an inner structured representation that consists of 1) a natural language question, 2) an inclusion-exclusion list that includes or excludes specific chemical descriptors for target catalysts and 3) a relational operator that describes how the search can be shifted from the previous query’s candidate catalysts to a different region in chemical space. We employ three relational operators: 1) similar-to, 2) subclass-of and 3) dissimilar-to. Each LLM answer represents a set of candidate catalysts, and each candidate is scored using a reward function. Starting the search with the root node, the search algorithm progresses by expanding each node into a set of children nodes with actions . Each layer of the search tree is pruned based on the reward of the nodes. Finally, when maximum search depth is achieved, we select the node with the highest reward as the overall answer to the initial prompt.
@@ 3.2 Planner-guided Search
Our planner component is responsible for systematically expanding the search by contextually determining viable actions. Specifically, the planner conditions action selection on the full sequence of previous queries and catalysts discoveries traced from the root of the search tree. This contextual grounding enables automatically constraining subsequent search directions in a scientifically coherent way.
Consider any node in the search tree where the planner generates the next query to execute (shown via orange box on top left in Figure 4). Next, the LLM executes the query and retrieves a set of candidate catalysts (such as Cu, Pd etc.). Each of these candidates are then transformed into a 3D atomistic representation and evaluated by a reward function described in sections 3.3 and 3.4. All candidates at a given depth in the search tree are collected and only a subset are chosen for expanding in the next iteration. The process repeats iteratively (shown via rows in above diagram) until the maximum tree depth is reached.
Overall, by leveraging language models to contextually expand search, ChemReasoner-Planner balances exploration with producing interpretable, scientifically-grounded reasoning paths. See section E for a complete trace of a planner guided search.
@@ 3.3 Reward via Structure Optimization and Energy Prediction
Each reward function returns a real number as a measure of the catalyst’s goodness (higher is better) for a given input question. In this work, we implement two reward functions with different levels of complexity.
Adsorption energy-based reward: This function returns the adsorption energy of the most stable binding configuration of the catalyst as the reward. The computation begins with translating the symbolic representation of the catalysts (such as “Platinum”) and adsorbates (e.g., “CO”) into a 3D atomistic structure (Fig. 4 right). The stability and energy of a catalyst’s atomic structure directly impacts its catalytic activity and selectivity. Therefore, we compute the most stable configuration for the catalyst-adsorbate pair and use its adsorption energy as a measure of the reward (Section A). The optimization process, also known as the relaxation process, iteratively relaxes the atomic positions of the 3D structure until an energy minimum is found. A GNN (Gasteiger et al., 2021) is then used to calculate the adsorption energy from this state.
Reaction pathway-based reward: This function measures the goodness of a catalyst considering multiple reaction pathways and intermediate steps. We initially obtained reaction pathways from the LLM. Given each reaction pathway, it computes the energy functions for every intermediate step. Figure 5 shows two instances of the same reaction pathway for two different catalysts. As the figure shows, proceeding from one reaction step to another requires different amounts of energy which depends on the catalyst. Intuitively, moving from a lower-energy state to higher-one can be viewed as a “hill-climbing” in the energy landscape (indicated by the red and blue arrow for two different catalysts) and we formulate a function that would assign the highest reward to pathways with the smallest hills to climb (Eq. 1). is the intermediate at step of the reaction and is energy of adsorption energy of on some catalyst. The overall reaction-based reward function for the top- catalysts is therefore,
@@ 4 Experiments
We conduct experiments to evaluate if a system using LLM-guided heuristic search with quantum-chemistry feedback can discover more novel and effective catalysts than state-of-the-art LLMs alone. Our experiments focus on three key research questions for enabling such a system.
@@ 4.1 Experimental setup
System Implementation LLMs in our experimental setup included OpenAI GPT-3.5 and GPT-4. Although we initially benchmarked with LLama2 (Touvron et al., 2023), we found that its instruction-following capabilities in this domain were too limited to allow evaluation. For our GNN reward model, the GemNet-dT model (Gasteiger et al., 2021) from the OpenCatalysis project was utilized. Runtime configurations and inference scaling performance for each of these models are provided in section C.6. Inferences for both OpenAI models were executed in parallel using asynchronous execution features. The GNN inferences were run on a single GPU on a DGX2/V100 and A100 systems.
@@ 5.1 Comparing Reasoning Approaches
These actions are sampled with uniform probability, without using the same criteria twice. On the contrary, ChemReasoner-Planner uses LLM-suggested actions for expanding the search space which do not require any expert specification.
@@ 5.2 Performance characterization of Key Components
ChemReasoner performs a significantly large number of LLM and GNN inferences that influence both its performance and throughput. The following factors control the computational complexity of the overall execution.
LLM Inference and Tree Search The runtime of ChemReasoner is , where is the number of nodes in the search tree. increases exponentially with the maximum depth (set to 5) and the branching factor for expanding each node in the search tree, denoted as (set to 8). We use a beam width (set to 6) parameter to control the number of nodes that are expanded in each iteration of the search. ChemReasoner performs LLM inferences for each node in the search tree to 1) plan actions, 2) suggest candidate catalysts, and 3) transform LLM answers into atomistic structures. It is important to note the strong intra-node and inter-node dependencies between these LLM queries, as the output from one query is fed into the next ones. This dependency exists both within a node and to its children in the search tree. Therefore, robustness of LLM instruction-following behavior, specifically the ability to return answers in a consistent format, becomes a critical factor towards an effective system. Overall, ChemReasoner-Expert and ChemReasoner-Planner executes and LLM inferences respectively.
Requirements for LLM Scaling and Robustness Therefore, to execute the entire benchmark of 145 queries with requires execution of 28,000-42,000 LLM inferences, which makes the the scalability and throughput of a generative LLM a critical factor for successful experimentation. By default, we batch LLM inferences in sizes of 48 and execute the calls asynchronously for scalability. These two factors, namely the robustness of instruction-following and high-throughput execution of batched queries restricted our experiments to only GPT-3.5 and GPT-4 models. Our experiments with LLama2 were both affected by frequent spurious answers as well as by the relatively low throughput on A100 systems. We recognize the importance of both development and benchmarking on open-source models and plan to pursue it as future work.
@@ 5.3 Considering DFT-Based Rewards
A critical component for a reasoning system is accurate feedback. For this reason, the use of DFT-based rewards would be the default choice. However, due to their significant computational cost, we elected to employ GNNs (Zitnick et al., 2020). We evaluated our top GNN results using DFT simulations; overall, we found that DFT calculation-based rewards diverge in some cases significantly from GNN predictions (Table 3). Here, we briefly examine the divergence between SOTA GNN models and ground truth DFT simulations.
One limitation of the GNN likely arises from the conversion of text-based catalyst recommendations from the LLM to 3D atomistic structures for the GNN calculation. Our structure generation method is limited to pre-defined lattice structures, (e.g., face centered cubic (FCC), body centered cubic (FCC), and hexagonal close packed (HCP)). While these lattice structures effectively describe mono-metallic structures, multimetallic compounds may exhibit more complex lattice structures. Thus, our assumption of simpler structures may be limiting. Furthermore, our multimetallic bulk structures were generated by randomly placing elements throughout the bulk, creating additional complexity.

### v4

@@ Abstract
The discovery of new catalysts is essential for the design of new and more efficient chemical processes in order to transition to a sustainable future. We introduce an AI-guided computational screening framework unifying linguistic reasoning with quantum-chemistry based feedback from 3D atomistic representations. Our approach formulates catalyst discovery as an uncertain environment where an agent actively searches for highly effective catalysts via the iterative combination of large language model (LLM)-derived hypotheses and atomistic graph neural network (GNN)-derived feedback. Identified catalysts in intermediate search steps undergo structural evaluation based on spatial orientation, reaction pathways, and stability. Scoring functions based on adsorption energies and reaction energy barriers steer the exploration in the LLM’s knowledge space toward energetically favorable, high-efficiency catalysts. We introduce planning methods that automatically guide the exploration without human input, providing competitive performance against expert-enumerated chemical descriptor-based implementations. By integrating language-guided reasoning with computational chemistry feedback, our work pioneers AI-accelerated, trustworthy catalyst discovery.111The code and datasets discussed in this paper are available at https://github.com/pnnl/chemreasoner.
@@ 1 Introduction
The discovery of new catalysts requires one to identify the optimal combination of chemical descriptors (or properties) and use these descriptors to propose catalysts. However, such descriptors are only empirically understood, presenting a challenge for computational studies of catalysis. Typically, chemists actively reason to mentally search through reactants, catalysts, and operating conditions that enable more energy-efficient chemical conversions. However, as discussed by Nørskov et al. (2011), linking microscopic surface properties to macroscopic catalytic performance via chemical descriptors remains a barrier to descriptor-based catalyst search.
Large Language Models (LLMs) (Wei et al., 2022; Ouyang et al., 2022; Taylor et al., 2022; Lai et al., 2023; OpenAI, 2023) offer a new opportunity to realize such a data-driven autonomous search to accelerate scientific discovery. Our work aims to enhance natural language reasoning capabilities with quantum-chemical feedback to discover optimal catalysts for target reactions.
The Challenge However, reasoning about complex catalytic processes requires modelling across multiple modalities, extending beyond the capabilities of existing language models. This includes merging scientific concepts from literature and property prediction with 3D atomistic configurations. Determining the best catalyst is a multi-step process, requiring reasoning about multiple macroscopic properties. The first step involves identifying an optimal set of chemical descriptors (e.g. “resistance to poisoning”, “porosity”) which are relevant to the reaction in question. Formally, given a set of these important descriptors , we want to identify the optimal subset to consider when suggesting new catalysts. This yields possible permutations to reason over. The amount of reasoning required scales combinatorially with the number of available properties, necessitating an autonomous reasoner: an LLM. Using its knowledge of scientific concepts, the LLM both proposes important properties and proposes the best possible catalysts (from all possible catalysts) that have these properties. Pruning this large space of candidate catalysts requires reasoning about the complex microscopic interactions that occur between atomistic structures in 3D space based on macroscopic properties (Figure 2). Further, while simple reactions can be assessed via adsorption energies of 3D chemical structures, complex reactions demand consideration of multi-step reaction pathways and selectivity (Unsleber et al., 2023).
@@
We demonstrate the decisive impact of planning methods in automatically navigating chemical search spaces over a SOTA LLM-based implementation. Our purely LLM-planned approach with zero human input (ChemReasoner-Planner) surpasses search guided by expert-selected chemical descriptors (ChemReasoner-Expert) for two out of three categories in our evaluation benchmark.
@@
Third, and uniquely, we establish the domain-grounding of language models via quantum chemical property feedback. We go beyond screening catalysts on adsorption energies alone and propose a methodology to reason in terms of reaction pathways and energy barriers.
Our work pioneers an AI-guided approach to computational catalyst screening and discovery. To facilitate community adoption and advancement of this extremely interdisciplinary and compute-intensive research, all of our datasets and code are freely available on github. This includes query benchmarks from catalysis experts, multi-modal provenance trails, over 700,000+ atomistic trajectories, and additional validation of catalyst candidates from density functional theory calculations.
@@ 2.2 LLMs for Chemistry
LLMs for chemistry can be divided into two categories: domain-specific models and adapted general-purpose models. Multimodal, domain-specific molecule-language models have recently emerged to target a number of problems in the chemistry domain (Edwards et al., 2021; Vall et al., 2021; Zeng et al., 2022; Xu & Wang, 2022; Su et al., 2022; Edwards et al., 2022a). For brevity, we discuss them further in Appendix D. On the other hand, general-domain LLMs such as GPT-4 (OpenAI, 2023) have been adopted in an agent-based approach to interface with chemistry-specific tools, allowing information gathering and hypothesis generation (Boiko et al., 2023; Bran et al., 2023). While this work is exciting, it differs from our approach, where we employ an overarching, domain-specific reward function to probe the LLM’s knowledge, enabling stronger, simulation-backed guarantees about complex scientific reasoning. We note that our method builds upon Sprueill et al. (2023). Unlike their approach, and the approach of Yao et al. (2023), which use LLM-computed rewards, we integrate true computational chemistry-based rewards to guide the model. Further, we integrate a context-aware planner into our search algorithm to automatically guide the search.
@@
.append()
3 System and Methods
3.1 Heuristic Search
The goal of our heuristic search is to answer a user-specified natural language query by systematically exploring candidates from different regions in the chemical space. Heuristic searches have been applied for general problem solving and reasoning tasks for LLMs (Yao et al., 2023). Our work follows the catalysis-focused approach of Sprueill et al. (2023), where the original query (or prompt) and corresponding LLM answers are systematically modified by applying different screening criteria to iteratively contextualize the LLM prompts and answers into a narrower region of the chemical space. This process is illustrated in Figure 2, where the addition of chemical descriptors change the catalysts that are recommended. Since the number of branching pathways could hypothetically expand exponentially, we employ a domain-specific reward function to prune candidate catalysts that don’t show promise of high catalytic activity.
More formally, our goal is to search through chemical descriptors and design constraints to determine the optimal prompt, which leads the LLM to return the best candidate catalysts for a catalysis related query. Starting with a general prompt , we use a set of actions to modify the prompt to improve the LLM output with respect to a reward function, . Notably, ChemReasoner-Planner generates its own action space (Figure 3).
Definitions: We define the Search Tree as a hierarchical tree consisting of (prompt, answer, and reward) nodes. Each node in this tree represents a state in the search space. Nodes are linked if an action modifies the prompt from one node to the next. We denote a path from root to leaf node a Reasoning Pathway.
Following Sprueill et al. (2023), each node contains a template LLM prompt and an internal structured representation that provides additional context to the template prompt. This internal representation consists of 1) a natural language question, 2) an inclusion-exclusion list that includes or excludes specific chemical descriptors for target catalysts and 3) a relational operator that describes how the search can be shifted from the previous query’s candidate catalysts to a different region in chemical space (i.e. similar-to or different-from the previous candidate catalysts). Starting the search with the root node, the search algorithm expands each node into a set of children nodes with a set of actions , which modify the internal representation of each node. The LLM then answers the modified prompt in each node, providing a set of candidate catalysts, and each candidate is scored using a reward function. Each layer of the search tree is pruned using a beam search algorithm a beam search algorithm (Rubin & Reddy, 1977), leaving only those nodes with the highest rewards. Finally, when the maximum search depth is reached, we select the node in the tree with the highest reward as the best answer to the initial prompt.
@@ 3.2 Planner-Guided Search
Our planner component is responsible for systematically expanding the search by contextually determining viable actions. Specifically, the planner selects actions based on the current context of the search, using vocabulary relevant to the current catalyst candidates. This contextual grounding automatically constrains the search direction in a scientifically coherent way. A technical description of the planner prompt is given in Appendix 6.
Consider any node in the search tree where the planner generates the actions for a given node (shown in the orange boxes of Figure 3). Next, we execute the actions, creating several child nodes, and retrieve a set of top- candidate catalysts from the LLM (such as Cu, Pd etc.). As shown in figure 4, each of these candidates are then transformed into a 3D atomistic representation and evaluated by a reward function, which approximates catalytic activity (see sections 3.3 and 3.4 for details). All nodes at a given depth in the search tree are collected and only a subset of nodes are chosen for further search in the next iteration, filtered by highest reward. The process repeats iteratively (shown via rows in above diagram) until the maximum tree depth is reached.
Overall, by leveraging the LLM to contextually expand the tree search, ChemReasoner-Planner explores candidate catalysts in a more interpretable, scientifically-grounded reasoning pathways. See section E for a complete trace of a planner guided search, using GPT-4.
@@ 3.3 Reward via Structure Optimization and Energy Prediction
Each reward function returns a real number as a measure of the catalyst’s goodness (higher is better) for a given input question. In this work, we implement two reward functions with different levels of complexity. The first reward function targets catalysts for the adsorption of particular chemical species, while the latter targets catalysts with higher approximate reactivity for certain reactions.
Adsorption Energy-Based Reward: To target catalysts for the adsorption of specific chemical species, we use an adsorption energy-based reward, which returns the adsorption energy of the most stable binding configuration of the catalyst as the reward. The computation begins with translating the symbolic representation of the catalysts (such as “Platinum”) and adsorbates (e.g., “*CO”) into a 3D atomistic structure (Fig. 4 right). The stability and energy of a catalyst’s atomic structure directly impacts its catalytic activity and selectivity. Therefore, we compute the most stable configuration for the catalyst-adsorbate pair and use its adsorption energy as a measure of the reward (Section A). The optimization process, also known as the relaxation process, iteratively relaxes the atomic positions of the 3D structure until an energy minimum is found. A GNN (Gasteiger et al., 2021) is then used to calculate the adsorption energy from this state.
[…]
Given each reaction pathway, our reward function computes the adsorption energies for every intermediate step. Figure 5 shows two instances of the same reaction pathway for two different catalysts. As the figure shows, proceeding from one reaction step to another requires different amounts of energy, which depends on the catalyst and adsorbate interaction. Intuitively, moving from a lower-energy state to higher-one can be viewed as a “hill-climbing” in the energy landscape (indicated by the red and blue arrow for two different catalysts) and we formulate a function that would assign the highest reward to pathways with the smallest hills to climb (Eq. 1). is the intermediate at step of the reaction and is the adsorption energy of on some catalyst. The overall reaction-based reward function for the top- catalysts is the average of , where in one of the top- catalysts and calculates the smallest maximum energy jump over all paths,
@@ 4 Experiments
We conduct an experiment to evaluate if ChemReasoner, using an LLM-guided heuristic search with quantum-chemistry feedback, can discover more novel and effective catalysts than state-of-the-art LLMs alone. Our experiments focus on three key research questions for enabling such a system.
@@ 4.1 Experimental setup
System Implementation The LLMs used in our experimental setup included OpenAI GPT-3.5 and GPT-4. Although we initially benchmarked ChemReasoner with LLama2 (Touvron et al., 2023), we found that its instruction-following capabilities in this domain were too limited to allow an evaluation.
For our GNN reward model, we utilized the GemNet-dT model (Gasteiger et al., 2021) from the Open Catalyst Project. Runtime configurations and inference scaling performance for each of these models are provided in section C.5. Inferences for both OpenAI models were executed in parallel using asynchronous execution features. The GNN inferences were run on a single DGX2/V100 or A100 GPU.
@@ 5.1 Reasoning Approaches
These actions are sampled with uniform probability, without using the same criteria twice. On the contrary, ChemReasoner-Planner uses LLM-suggested actions for expanding the search space which do not require any expert specification. As such, ChemReasoner-Planner’s actions can adapt to the changing context of the search, whereas ChemReasoner-Expert’s action space remains static.
@@ 5.2 Performance Characterization of Key Components
ChemReasoner performs a large number of LLM and GNN inferences that influence both its performance and throughput. The following factors control the computational complexity of the overall execution.
LLM Inference and Tree Search The runtime of ChemReasoner is , where is the number of nodes in the search tree. increases exponentially with the maximum depth (set to 5) and the branching factor for expanding each node in the search tree, denoted as (set to 8). We use a beam width parameter (set to 6) to control the number of nodes that are expanded in each iteration of the search. ChemReasoner performs LLM inferences for each node in the search tree to 1) plan actions, 2) suggest candidate catalysts, and 3) transform LLM answers into atomistic structures. It is important to note the strong intra-node and inter-node dependencies between these LLM queries, as the output from one query is fed into the next ones. This dependency exists both within a node and to its children in the search tree. Therefore, robustness of LLM instruction-following behavior, specifically the ability to return answers in a consistent format, is a critical factor for an effective system. Overall, ChemReasoner-Expert and ChemReasoner-Planner executes and LLM inferences respectively.
To execute the entire benchmark of 145 queries with requires execution of 28,000-42,000 LLM inferences, which makes the the scalability and throughput of a generative LLM a critical factor for successful experimentation. By default, we batch LLM inferences in sizes of 48 and execute these inferences asynchronously for scalability. These two factors, namely the robustness of instruction-following and high-throughput execution of batched queries restricted our experiments to only GPT-3.5 and GPT-4 models. Our experiments with LLama2 were both affected by frequent spurious answers and relatively low throughput on A100 systems. We recognize the importance of the development and benchmarking of open-source models, but this particular topic should be the subject of future work.
@@ 5.3 Considering DFT-Based Rewards
A critical component for a reasoning system is accurate feedback. For this reason, the use of DFT-based rewards would be the gold standard reward calculations. However, due to their significant computational cost, we are forced to employ GNNs for high-throughput evaluation of catalyst candidates (Zitnick et al., 2020). We evaluated our top GNN results using DFT simulations; overall, we found that DFT calculation-based rewards diverge in some cases significantly from GNN predictions (Table 3). Here, we briefly examine the divergence between SOTA GNN models and ground truth DFT simulations.
One limitation of the GNN likely arises from the conversion of text-based catalyst recommendations from the LLM to 3D atomistic structures for the GNN calculation. Our structure generation method (see Section A) is limited to pre-defined lattice structures, (e.g., face centered cubic (FCC), body centered cubic (FCC), and hexagonal close packed (HCP)). While these lattice structures effectively describe mono-metallic structures, multimetallic compounds may exhibit more complex lattice structures. Thus, our assumptions of reference structures may not adequately reflect realistic catalyst structures. Furthermore, our multimetallic bulk structures were generated by randomly placing elements throughout the bulk, creating additional complexity due to the emergence of defects in the bulk structures.